On January 1, 2018, Algerian Delivery had 100,000 shares of common stock outstanding. The following transactions occurred during 2018: March 1: Reacquired 3,800 shares, accounted for as treasury stock. September 30: Sold all the treasury shares. December 1: Sold 12,800 new shares for cash. December 31: Reported a net income of $305,750.

Required: Calculate Algerian Delivery's basic earnings per share for the year ended December 31, 2018. (Round your answer to 2 decimal places.)
Answer #1

Ans:

Net Income = 305,750

Calculation of Weighted Average of Number of Equity Shares:

Outstanding on Jan 1 = 100,000*12/12 = 100,000

Less: Treasury Stock 3800*10/12 = (3167)

Sold Treasury Shares 3800*3/12 = 950

New Shares Sold 12,800*1/12 = 1067

Weighted Average of Equity Shares during the year = 98,850

Basic Earnings per Share = Net Income/Weighted Average of Equity Shares outstanding during the year

= 305,750/98,850 = $3.09